# RIKEN-iTHEMS-Report-25, KYUSHU-HET-308

### v2

@@ 3 Conclusion
The pattern which the ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ exhibits as the function of the flux, such as that in Figs. 2 and 3, clearly indicates the quantum phase of the system tHooft:1979rtg ; tHooft:1981sps . In this paper, by employing the halfway HMC Abe:2024fpt for the simplest example the $SU(2)$ pure Yang–Mills theory, we illustrated a direct Monte Carlo calculation of the pattern is feasible at least as far as the integrand of the functional integral is real-positive; when the complex phase of the integrand is small, it could be treated by reweighting. Our methodology itself is quite general so we hope to carry out analyses in more intriguing situations: It should be interesting to study the finite temperature cases (such as in Ref. deForcrand:2001nd ) to observe the confining/deconfining phase transition. The inclusion of matter fields such as the adjoint Higgs and adjoint fermions will give a more intricate pattern of ’t Hooft partition function. We hope to attack these problems in near future.
@@ Acknowledgments
We would like to thank Motokazu Abe for discussions spanning over a long period of time which finally led to this work. We also thank Philippe de Forcrand, Yui Hayashi, Yuki Nagai, Soma Onoda, Yuya Tanizaki, Akio Tomiya, and Hiromasa Watanabe for helpful discussions. We appreciate the opportunity of the discussion during the YITP–RIKEN iTHEMS conference “Generalized symmetries in QFT 2024” (YITP-W-24-15) in execution of this work. Numerical computations in this paper were carried out on Genkai, a supercomputer system of the Research Institute for Information Technology (RIIT), Kyushu University. O.M. acknowledges the RIKEN Special Postdoctoral Researcher Program. The work of H.S. was partially supported by Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research, JP23K03418.
@@ Appendix A Euclidean $90^{\circ}$ discrete rotations on the $\mathbb{Z}_{2}$ flux in 4D and 3D
On components of a second rank anti-symmetric tensor,
@@
Euclidean $90^{\circ}$ rotations in a $\mu\nu$-plane are represented as $B_{a}\to\sum_{b=1}^{6}(\Lambda_{\mu\nu})_{ab}B_{b}$, where the representation matrices are given by,
@@
For finite temperature, $\mathcal{Z}[B]$ is invariant only under 3D spatial $90^{\circ}$ rotations generated by $\Lambda_{12}$, $\Lambda_{13}$, and $\Lambda_{23}$. Then each irreducible representation is decomposed into smaller multiplets labeled by the roman index in the equation number, such as “a” in (A.5a). We may then take the average of $\mathcal{Z}[B]$ over elements within these smaller multiplets.

### v3

@@ 3 Conclusion
The pattern exhibited by the ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ as the function of the flux, such as that in Figs. 2 and 3, clearly indicates the quantum phase of the system tHooft:1979rtg ; tHooft:1981sps . In this paper, by employing the halfway HMC Abe:2024fpt for the simplest example, the $SU(2)$ pure Yang–Mills theory, we illustrated that a direct Monte Carlo calculation of the pattern is feasible, at least as long as the integrand of the functional integral is real-positive; when the complex phase of the integrand is small, it could be treated by reweighting. Our methodology itself is quite general so we hope to carry out analyses in more intriguing situations: It should be interesting to study finite temperature cases (such as in Ref. deForcrand:2001nd ) to observe the confining/deconfining phase transition. The inclusion of matter fields such as the adjoint Higgs and adjoint fermions will give a more intricate pattern of the ’t Hooft partition function. We hope to tackle these problems in the near future.
@@ Acknowledgments
We would like to thank Motokazu Abe for discussions spanning a long period of time, which finally led to this work. We also thank Philippe de Forcrand, Yui Hayashi, Yuki Nagai, Soma Onoda, Yuya Tanizaki, Akio Tomiya, and Hiromasa Watanabe for helpful discussions. We appreciate the opportunity of discussion during the YITP–RIKEN iTHEMS conference “Generalized symmetries in QFT 2024” (YITP-W-24-15) in the execution of this work. The numerical computations in this paper were carried out on Genkai, a supercomputer system of the Research Institute for Information Technology (RIIT), Kyushu University. O.M. acknowledges the RIKEN Special Postdoctoral Researcher Program. The work of H.S. was partially supported by a Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research, JP23K03418.
@@ Appendix A Euclidean $90^{\circ}$ discrete rotations on the $\mathbb{Z}_{2}$ flux in 4D and 3D
On the components of a second rank antisymmetric tensor,
@@
Euclidean $90^{\circ}$ rotations on the $\mu\nu$-plane are represented as $B_{a}\to\sum_{b=1}^{6}(\Lambda_{\mu\nu})_{ab}B_{b}$, where the representation matrices are given by,
@@
For finite temperatures, $\mathcal{Z}[B]$ is invariant only under 3D spatial $90^{\circ}$ rotations generated by $\Lambda_{12}$, $\Lambda_{13}$, and $\Lambda_{23}$. Then each irreducible representation is decomposed into smaller multiplets labeled by a roman index in the equation number, such as “a” in (A.5a). We may then take the average of $\mathcal{Z}[B]$ over elements within these smaller multiplets.